OpenStudy (anonymous):

@Demonx341 can u help me ?

OpenStudy (anonymous):

question?

OpenStudy (anonymous):

sorry no

OpenStudy (anonymous):

no what ?

OpenStudy (anonymous):

let me see if anyone i know is good at this

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

everyone that is good at math that i have fanned are offline xD

OpenStudy (anonymous):

fine then thx anyway but i have more easy math other than this can u help then ?

OpenStudy (anonymous):

let me see it

OpenStudy (anonymous):

i just hate grids

OpenStudy (anonymous):

my guess is B

OpenStudy (anonymous):

nope Try again can get rid of D and B

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

no d and b are incorrect so pick form c or a

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Correct :D

OpenStudy (anonymous):

because You add 8+8+3+3 to go perimater

OpenStudy (anonymous):

can u help with another one please

OpenStudy (anonymous):

8*2 = 16 and 3*2 = 6

OpenStudy (anonymous):

bvrb

OpenStudy (anonymous):

what

OpenStudy (anonymous):

afk for little

OpenStudy (anonymous):

what does afk mean

OpenStudy (anonymous):

away from keyboard

OpenStudy (anonymous):

so r u back or not

OpenStudy (anonymous):

not yet

OpenStudy (anonymous):

but ur typing now?

OpenStudy (anonymous):

<--- his brother

OpenStudy (anonymous):

he had to do something

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

can u help me then his brother

OpenStudy (anonymous):

i can try

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

let me look

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do 5*10 = ammount of ft btw i suck at math so xD

OpenStudy (anonymous):

<--- still demon's brother

OpenStudy (anonymous):

50 and i know

OpenStudy (anonymous):

next? or last one?

OpenStudy (anonymous):

so is that teh answer

OpenStudy (anonymous):

50

OpenStudy (anonymous):

u said u suck at math so i dont want to fail lol

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

@DivineSolar

OpenStudy (anonymous):

there i hope he can help

OpenStudy (anonymous):

he does thx

OpenStudy (anonymous):

@undeadknight26

OpenStudy (anonymous):

Divine said he is doing a test

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

if you would like you can close this and say i need someone Great at math in the question and add fan/medal to it and people will help you

OpenStudy (anonymous):

or do that xD